# Evaluating Sample Utility for Data Selection by Mimicking Model Weights

### v1

@@ Abstract
Foundation models rely on large-scale web-crawled datasets, which frequently contain noisy data, biases, and irrelevant content. Existing data selection techniques typically use human heuristics, downstream evaluation datasets, or specialized scoring models, and can overlook samples’ utility in the training process. Instead, we propose a new approach, Mimic Score, a data quality metric that uses a pretrained reference model as a guide to assess the usefulness of data samples for training a new model. It relies on the alignment between the gradient of the new model parameters and the vector pointing toward the reference model in weight space. Samples that misalign with this direction are considered low-value and can be filtered out. Motivated by the Mimic score, we develop Grad-Mimic, a data selection framework that identifies and prioritizes useful samples, automating the selection process to create effective filters. Empirically, using Mimic scores to guide model training results in consistent performance gains across six image datasets and enhances the performance of CLIP models. Moreover, Mimic scores and their associated filters improve upon existing filtering methods and offer accurate estimation of dataset quality.
@@ 1 Introduction
Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], LLaMa-3.1 [3], and GPT-4 [4]. These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content from their web sources. To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the model development pipeline [5, 6]. For example, the FineWeb dataset [7], containing 15 trillion tokens used for training large language models, is created through eight carefully designed filtering steps meant to refine raw web content, ensuring high model performance.
While effective, such filtering recipes face several limitations. The choices of filtering steps and rules often rely on handcrafted heuristics and require domain expertise and expensive experimentation. Such filters often fail to provide fine-grained insights into individual samples and overlook samples’ utility in training. These lead to coarse selection and suboptimal model performance.
Existing techniques often suggest selecting samples based on semantic similarity to downstream evaluation datasets [8, 9], using specialized filtering networks [10, 11, 12, 13], or training influence models to score samples [14, 15]. Yet, these approaches require access to additional datasets or involve specialized model training, adding dependencies and complexity.
We introduce the Mimic Score, a new data quality metric that assesses sample contribution in the weight updates using a pretrained reference model. We show that by leveraging a model located in a more optimal part of the weight space, we bypass the need for downstream datasets or specialized training, and can use it as a selection guide. Our approach is based on the alignment between each sample’s gradient and the direction towards reference model in weight space.
Samples that potentially pull the model in undesirable directions—misguiding weight updates—are considered low-value and should be filtered out.
@@
  • Stage 1: During training, Grad-Mimic prioritizes samples to learn by using mimic score, steering the model toward the reference model’s weight space.
  • Stage 2: After training, Grad-Mimic identifies sample utility across training steps and aggregates these assessments for an ensemble filter, automating data selection.
We validate the effectiveness of Grad-Mimic through various experimental setups, presenting empirical results at each stage. First, we create a synthetic and controlled scenario that adds different levels of label noise. We demonstrate its capability to accurately identify mislabeled samples and provide a reliable estimation of overall dataset quality. Compared to the standard training method, Grad-Mimic successfully down-weights the contribution of undesirable samples and enhances model performance across six image datasets.
Next, we test Grad-Mimic in more challenging settings: using large-scale web-crawled multimodal data. We train CLIP models [2] from scratch on 10 million and 100 million samples from the DataComp dataset [16]. We use publicly available CLIP model weights as our reference and demonstrate that mimic scores help navigate training, resulting in performance gains across both scales. Compared to human-designed filters, our refined dataset achieves higher model performance. Furthermore, mimic scores complement CLIP score-based filters [17] by removing low-value samples, lifting model performance with less data.
@@
  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between per-sample gradient and the vector toward the reference model’s weights.
  • Automated Data Selection Framework: Building on the mimic score, we present Grad-Mimic, a framework that effectively identifies useful samples to improve training and automates effective data selection.
  • Dataset Quality Assessment: The mimic score, along with its developed filters, improves existing filtering strategies and provides reliable estimations of dataset quality.
@@ 2 Related Work
Figure 1: High-value vs. Low-value Samples Identified by Mimic Score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their mimic scores. Below each image, we show their caption and corresponding CLIP score (where a higher score suggests good quality [17]). Mimic scores closely align with their CLIP scores. High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
@@
Data selection techniques can generally be categorized into group- and sample-level approaches. Group-level approaches focus on optimizing the mixture of data domains to curate high-quality datasets [18, 19, 20, 21]. Sample-level methods, which are the focus of Grad-Mimic, aim to filter out individual noisy and unhelpful samples. Prior research has explored various strategies, including identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training update [24, 25]. These methods often rely on batch-level gradients, which can struggle with large amounts of noise in a mini-batch. Grad-Mimic addresses this challenge with the help of a reference model, acting as a reliable guide to select samples.
@@
Multimodal models are typically trained on large-scale web datasets [26, 27, 16, 28], which are often noisy and require careful curation. Previous approaches have included selecting samples based on their semantic similarity to the downstream evaluation datasets [8, 9], using semantic deduplication [29], developing specialized filtering networks [10, 11, 12, 13], or training influence models [14, 15]. While effective, they require access to target datasets or introduce additional training complexities. Grad-Mimic overcomes these limitations and offers a more efficient alternative: using pretrained weights to identify useful samples.
@@
Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [30, 31, 32, 33, 34]. These estimates typically come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [35, 36], often encoded as labeling functions. The outputs of these labeling functions are modeled to assess their reliability and then aggregated to produce final labels [30]. Weak supervision has demonstrated success in various domains, including segmentation tasks [37], relation extraction [38], MRI sequence analysis [39], sensor data [40], chemical reaction extraction [41], and enhancing predictions from pretrained models [42, 43]. Unlike most existing works focusing on label aggregation for data annotation, Grad-Mimic uses weak supervision for filtering purposes. We mitigate assessment noise across training steps, aggregating filter outputs into an effective ensemble filter.
@@ 3 Evaluating Sample Utility
Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients. Then it adapts alignment scores to reweight gradients that contributed to the weight movements. In Stage 2, derived alignment scores, mimic scores, are used to identify low-value samples and build an ensemble filter, enabling accurate noise detection, precise dataset quality assessment, and training dataset discovery.
We quantify samples by their contributions to the learning process. Our principle is that samples that potentially pull the model in undesirable directions, thereby misdirecting weight updates, should be considered low-value.
We start with setup and notation, then explain how our scoring metric is derived and used in our data selection framework, Grad-Mimic.
@@ 3.1 Mimic Score Calculation
To evaluate whether a sample steers the model in an undesirable direction, we use a pretrained model as a reference. This reference model, denoted by $\theta_{\text{ref}}$, resides in a more optima part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$. We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample’s utility in approximating a better weight configuration. These reference weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [44, 45].
The reference model can be obtained in two ways: either by training it on the dataset to achieve a target performance, if resources permit, or more efficiently, by using publicly available pretrained models, which eliminate the need for data access and expensive training.
The vector pointing toward the reference model, at training step $t$, is represented by $v_{t}:=\theta_{\text{ref}}-\theta_{t}$. We examine how each sample’s negative gradient $-g_{i,t}$, intended for updating model weights, aligns with vector $v_{t}$. We measure the alignment degree by considering both the direction and magnitude of the negative gradient. Specifically, we compute the projection length of $-g_{i,t}$ onto $v_{t}$, yielding an alignment score $m_{i,t}$, computed as follows
@@ 4 Grad-Mimic: Data Selection Framework
The mimic score represents each sample’s usefulness in guiding the model in a better direction. Building on this metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes samples to learn and then identifies useful samples to design an effective filter. A complete Grad-Mimic workflow is illustrated in Fig. 2.
@@ 4.2 Stage 2: Automated Filter Design
The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, automate data selection based on these assessments, and estimate dataset quality.
@@ 4.2.1 Sample Identification
  • Threshold-based Selection: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
  • 1D Clustering: Samples are categorized into two groups using clustering techniques such as $k$-means clustering [46] or Gaussian Mixture Models (GMMs), allowing more automated selection based on their assigned cluster.
  • Top-k Percent Selection: Samples are ranked by mimic score, and the top-$k$ percent are chosen. The value of $k$ can be adjusted based on the available training budget.
After identification, we binarize mimic scores and categorize samples into “retain” or “discard” at each training step. We treat these assigned groups as filter outputs and then combine them for a final filter.
@@ 4.2.2 Filter Output Aggregation
While filtering at an individual training step can refine the dataset, relying on isolated filters risks overlooking training dynamics. Additionally, considering sample utility across different stages may yield complementary signals.
To address this, Grad-Mimic leverages weak supervision techniques [30, 31, 32, 33, 34], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps. We begin with learning a model to evaluate the reliability of each step assessment [32]. Once established, this model aggregates filtering decisions into a high-quality ensemble filter. Through aggregation, Grad-Mimic mitigates assessment noise and captures training dynamics. Grad-Mimic employs the Snorkel framework [31], a widely used method in the weak supervision community.
Ultimately, the ensemble filter selects high-value samples for a refined dataset. We can estimate dataset quality by analyzing the proportion of retained samples or using samples’ mimic scores as an overall metric. We summarize steps in Grad-Mimic in Algorithm 1 placed in Appendix B.
@@ 5 Experiments
We assess the effectiveness of mimic score within Grad-Mimic using two experimental setups across datasets of varying scales and domains. We first test in a controllable setting, adding noise into sample labels (Sec. 5.1), followed by evaluations on large-scale web-crawled datasets (Sec. 5.2). Our goals are to validate the following claims in both setups:
  • Enhanced Model Performance: The reference model acts as a reliable guide. Training a new model with samples prioritized by mimic scores enhances performance.
  • Accurate Sample Identification: Grad-Mimic effectively detects noisy samples, automates data selection, and complements existing filtering methods.
  • Dataset Quality Assessment: Mimic scores and their associated filters offer reliable estimates of dataset quality, highly correlated with noise levels and performance gains.
5.1 Simulating Mislabeled Samples
Setups.
We begin with a controlled experiment by adding various levels of label noise to six image classification datasets. They are DTD [47], Flowers102 [48], STL10 [49], OxfordIIIT Pet [50], CIFAR10 [51], and CIFAR100 [51]. We fine-tune ViT-B/16 models [52] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples. We normalize mimic scores with a temperature of 0.5 and use a batch size of 32. We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last layer weights as the reference to navigate training on the noisy datasets. We detail more training configurations in Appendix C.1.
After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and clustering methods using $k$-means and GMM, then aggregate filter outputs across training steps using Snorkel framework [31].
Expected Results.
We aim to validate that a pretrained reference model can serve as a strong guide for model training. Moreover, we expect that mimic scores obtained from Grad-Mimic can help identify low-value, mislabeled samples.
@@ 5.1 Simulating Mislabeled Samples
Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, leading to improved denoising and performance gains.
@@
Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score metric. Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and identification methods. Notably, its performance remains high and robust regardless of the noise level.
@@
We compare Grad-Mimic’s training method with other approaches that use gradient information for sample prioritization. We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch. (ii) GraNd [22] reweights samples based on their gradient norm, prioritizing data that induce greater changes. (iii) AGRA [23] computes the cosine similarity of each sample gradient to an average gradient of a random batch, then excludes outlier samples. (iv) Grad-Match [24] adapts gradients by solving a subset selection problem to identify key samples within each batch.
@@
We present Grad-Mimic’s two-stage results separately. The Stage 1 results, shown in Table 1, report testing accuracy across different methods and noisy datasets using full fine-tuning and linear probing. We evaluate Grad-Mimic’s robustness by averaging accuracy across datasets under different levels of label noise. Results show that Grad-Mimic effectively identifies and down-weights mislabeled samples, improving denoising and closing the performance gap to the reference model. Notably, in the full fine-tuning setup where we only target the last layer to mimic, we successfully guide the entire model weight updates, ultimately outperforming other methods. This also highlights the computational feasibility of our approach. These findings validate the idea that using the direction toward a better-performing model can guide model training and lead to enhanced performance.

### v2

@@ Abstract
Foundation models are trained on large-scale web-crawled datasets, which often contain noise, biases, and irrelevant information. This motivates the use of data selection techniques, which can be divided into model-free variants—relying on heuristic rules and downstream datasets—and model-based, e.g., using influence functions. The former can be expensive to design and risk introducing unwanted dependencies, while the latter are often computationally prohibitive. Instead, we propose an efficient, model-based approach using the Mimic Score, a new data quality metric that leverages the weights of a reference model to assess the usefulness of individual samples for training a new model. It relies on the alignment between gradients and a target direction induced by the reference model. Using the derived Mimic Scores, we develop Grad-Mimic, a framework that prioritizes samples for learning, creates effective filters, and automates data selection. Empirically, using Mimic Scores to guide training improves data efficiency, results in consistent performance gains across six image datasets, and includes enhancements to CLIP models. Moreover, Mimic Score-based filters improve upon existing filtering methods, e.g., cutting 4.7 million samples to train better CLIP models while offering accurate estimation of training dataset quality.
@@ 1 Introduction
Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], and GPT-4 [3]. These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content. To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the development pipeline [4, 5]. For example, the FineWeb dataset [6], with 15 trillion tokens, is refined through eight carefully designed filtering steps to ensure high model performance.
Existing data selection methods have several limitations. We can divide these into two categories: those that are independent of a model being trained (or using pretrained models) and those that rely on such a model. Model-free approaches, such as those based on handcrafted heuristics or semantic similarity to downstream datasets [7, 8] often require expensive experimentation to establish selection rules and add unwanted dependencies on additional datasets. Furthermore, these methods can fail to provide fine-grained insights into individual samples and overlook sample utility in training, resulting in suboptimal performance. In contrast, model-based techniques, e.g., training specialized filtering networks  [9, 10, 11, 12] or using data influence functions  [13, 14], introduce complexity and are often computationally expensive.
Figure 1: High-value vs. low-value samples identified by the Mimic score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their Mimic scores. Each image is accompanied by its caption and corresponding CLIP score (higher scores suggest better quality). Mimic scores align closely with CLIP scores. High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
To address these limitations, we introduce Mimic Score, a new data quality metric that assesses sample’s contribution to weight updates using reference model’s weights. This is an efficient model-based technique. We show that by leveraging pretrained model weights (located in a more optimal part of the weight space), we bypass the need for downstream datasets or specialized model training and can use it as our selection guide. Our approach is based on the alignment between each sample’s negative gradient and a direction pointing toward the reference model in the weight space.
We argue that samples that potentially pull the model in undesirable directions—thereby misguiding weight updates—should be considered low-value and ruled out.
Leveraging model weights to evaluate sample utility offers several advantages. First, with the growing availability of well-trained models, adopting model weights has become more accessible than ever. Second, unlike other reference model-based approaches that require loading a second model and performing full inference passes [15, 16], our method scores each sample directly from the weights, improving efficiency.
@@
  • Stage 1: During training, Grad-Mimic prioritizes which samples to learn based on their mimic scores, improving data efficiency while steering the model toward the reference model’s weight space.
  • Stage 2: After training, Grad-Mimic uses mimic scores to evaluate each sample’s utility across training steps. These assessments are then aggregated to create an ensemble filter, which can further refine the dataset.
We validate Grad-Mimic through diverse experiments. First, in a controlled setting with varying levels of label noise, Grad-Mimic accurately identifies mislabeled samples and estimates training dataset quality. During training, it effectively down-weights undesirable samples and improves data efficiency across six image datasets. Next, we test it on large-scale web-crawled multimodal data by training CLIP models from scratch [2] on 10M and 100M samples from DataComp  [17]. Using publicly available pretrained weights as our reference, mimic scores effectively navigate training. The resulting filter performs better than human-designed ones and can complement CLIP score-based filters [18], lifting CLIP model performance with less training data.
@@
  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between gradients and the vector toward the reference model in the weight space.
  • Automated Data Selection Framework: Using the mimic score, we present Grad-Mimic, a framework that effectively identifizes useful samples to improve data efficiency and automates effective data selection.
  • Dataset Quality Assessment: The mimic score, along with its resulting filters, improves existing filtering strategies and provides reliable estimation of dataset quality.
@@ 2 Related Work
Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients. Then it adapts alignment scores to reweight gradients that contributed to the weight updates. In Stage 2, derived alignment scores—mimic scores—are used to identify low-value samples and build an ensemble filter, which can be used for noise detection, dataset quality estimation, and training dataset discovery.
@@
Data selection techniques can be categorized into group- and sample-level approaches. Group-level approaches focus on optimizing data domain mixtures to curate high-quality datasets [19, 20, 21]. Sample-level methods, which are the focus of Grad-Mimic, aim to provide a score measuring sample utility and filter out unhelpful samples. Prior research has explored various gradient-based strategies, such as identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training [24, 25]. These methods can struggle with large amounts of noise in batch gradients. Grad-Mimic overcome this with the help of a reference model.
@@
Multimodal models rely on large-scale web datasets [26, 17, 27], which are often naturally noisy and require careful curation. Previous approaches have included selecting samples based on semantic similarity to the downstream evaluation datasets [7, 8], using semantic deduplication [28], developing specialized filtering networks [9, 10, 11, 12], and using data influence functions [13, 14]. While effective, they require access to target datasets, introduce training complexities, and can be expensive in computation. Grad-Mimic overcomes these and offers a more efficient alternative: using public available model weights to identify useful samples.
@@
Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [29, 30, 31, 32]. These estimates can come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [33, 34], often encoded as labeling functions. The outputs of these labeling functions are modeled and then aggregated to produce final labels [29]. Weak supervision has demonstrated success across diverse domains [35, 36, 37, 38, 39]. Most prior works focus on weak label aggregation to construct datasets, Grad-Mimic leverages weak supervision for a new purpose: modeling sample utility assessments for a combined filtering decision.
@@ 3 Evaluating Sample Utility
We quantify samples by their contributions to the learning process. Our principle is that samples that potentially pull the model in an undesirable direction, thereby misdirecting weight updates, should be considered low-value.
We start with setup and notation, then explain how the scoring metric is derived and used in our data selection framework, Grad-Mimic.
@@ 3.1 Mimic Score Calculation
To evaluate whether a sample’s gradient pulls the model in an undesirable direction, we use the weights of a pretrained reference model as our guide. These reference weights, denoted by $\theta_{\text{ref}}$, reside in a more optimal part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$. We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample utility in approximating a better weight configuration ($\theta_{\text{ref}}$). These weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [40, 41].
Reference weights can be obtained in two ways: by training a model on the dataset to achieve a desired performance level, if resources permit, or more efficiently, by using publicly available well-trained models, which are now more accessible and can eliminate the need for direct data access or possible costly training.
The vector pointing toward reference model’s weights, at training step $t$, is represented by $v_{t}:=\theta_{\text{ref}}-\theta_{t}$. We examine how each sample’s negative gradient $-g_{i,t}$, intended for updating model weights, aligns with vector $v_{t}$. We measure this alignment degree by considering both the direction and magnitude of the negative gradient. Specifically, we compute the projection length of $-g_{i,t}$ onto $v_{t}$, yielding an alignment score $m_{i,t}$, computed as follows
@@ 4 Grad-Mimic: Data Selection Framework
Building on this simple but effective metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes which samples to learn and then identifies useful samples to design an effective filter. The complete Grad-Mimic workflow is illustrated in Fig. 2.
@@ 4.2 Stage 2: Automated Filter Design
The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, creating a combined filter based on sample utility assessments.
@@ 4.2.1 Sample Identification
  • Threshold-based: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
  • 1D Clustering: Samples are categorized into two groups using clustering such as $k$-means clustering [42] or Gaussian Mixture Models (GMMs), allowing more automated identification based on their assigned cluster.
  • Top-k Percent: Samples are ranked by their mimic score, and the top-$k$ percent are chosen. The value of $k$ can be adjusted based on the available training budget.
After identification, normalized mimic scores are binarized, and samples are categorized into “retain” or “discard” at each training step. We treat these assigned groups as each training step’s votes and then combine for a final filter.
@@ 4.2.2 Filter Output Aggregation
Although using the assessment at a training step can refine the dataset, relying on such isolated filters risks overlooking training dynamics. Moreover, evaluating a sample’s utility across all training stages may yield complementary signals.
To address this, Grad-Mimic uses weak supervision techniques [29, 30, 31, 32], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps. We start by learning a model to evaluate the reliability of each step’s assessment [31]. Once established, this model aggregates filtering decisions into a high-quality ensemble filter. Through aggregation, Grad-Mimic mitigates assessment noise and captures training dynamics. We employ the Snorkel framework [30], a widely used method in the weak supervision community.
Ultimately, we can use this ensemble filter to select high-value samples for a refined dataset. Moreover, we can use it to estimate training dataset quality by either analyzing the proportion of retained samples or by averaging samples’ mimic scores as an overall metric. We summarize the steps of Grad-Mimic in Algorithm 1 and detailed in Appendix A.
@@ 5 Experiments
We evaluate the effectiveness of Grad-Mimic using two experimental setups across datasets of varying scales and domains. We first test in a controllable setting, adding noise into sample labels (Sec. 5.1), followed by evaluations on large-scale web-crawled datasets (Sec. 5.2). Our goals are to validate the following claims in both setups:
  • Enhanced Data Efficiency: A pretrained reference model serves as a reliable guide, enabling mimic scores to prioritize samples that improve model performance.
  • Accurate Sample Identification: Grad-Mimic can detect noisy samples, construct effective filters, and automate data selection.
  • Dataset Quality Assessment: Mimic scores and their resulting filters provide reliable estimates of dataset quality.
@@
Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, improving data efficiency and closing the performance gap to reference model.
@@
Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score metric. Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and proposed identification methods. Notably, its performance remains high and robust regardless of the noise level.
5.1 Simulating Mislabeled Samples
Setups.
We begin with a controlled experiment by adding various levels of label noise to six image classification datasets. They are DTD [43], Flowers102 [44], STL10 [45], OxfordIIIT Pet [46], CIFAR10, and CIFAR100 [47]. We fine-tune the ViT-B/16 model [48] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples. We normalize mimic scores with a temperature of 0.5 and use a batch size of 32. We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last-layer weights as our reference to navigate training on the noisy datasets. We detail more training configurations in Appendix B.1.
After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and using $k$-means and GMM to cluster. Then we aggregate filter outputs across training steps using Snorkel framework [30].
@@ 5.1 Simulating Mislabeled Samples
Expected Results.
We aim to validate that a pretrained reference model’s weights can serve as a reliable guide. Moreover, we expect that mimic scores can help identify low-value, mislabeled samples.
@@
We compare Grad-Mimic’s training method (Stage 1) with other approaches that use gradient information for sample prioritization. We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch. (ii) GraNd [22] reweights samples based on their gradient norm, prioritizing data that induce greater changes. (iii) AGRA [23] computes the cosine similarity of each sample gradient to an average gradient of another random batch, then excludes outlier samples. (iv) Grad-Match [24] adapts gradients by solving a subset selection problem to identify key samples within each batch. Furthermore, we compare Grad-Mimic with Rho-Loss [15], a reference model-based approach that prioritizes training on worth learning samples, as well as influence function-based methods. Results are presented in Appendix C. Grad-Mimic outperforms both techniques while additionally offering efficiency advantages.
@@
We present Grad-Mimic’s two-stage results separately. Stage 1 results, shown in Table 1, report testing accuracy across various methods and noisy datasets using both training configurations. These results demonstrate that Grad-Mimic effectively identifies and down-weights mislabeled samples during training, enhancing denoising and narrowing the performance gap with the reference model. Notably, in the full fine-tuning setup—where we only target the last layer to mimic—Grad-Mimic successfully guides weight updates for the entire model, ultimately outperforming other methods. This highlights its computational efficiency, as it avoids requiring sample gradients from the full set of model weights. Additionally, we explore the impact of reference model quality in Appendix C, further strengthening our approach. These findings validate the idea that using the direction induced by a better-performing model can guide model training and improve data efficiency.
Stage 2 Results.
Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples. We use F1-scores to report the detection results in Table 2. Grad-Mimic accurately identifies mislabeled data across datasets and remains robust when facing different noise levels.
We use CIFAR100 dataset as an example and visualize computed mimic score distribution at each training step in Figure 3. In this dataset, half of the labels are flipped. We can see their distributions clearly separate into two groups, confirming that mimic scores are informative signals to identify samples. Despite variations over training, these dynamics can be captured in our aggregation step.
We use the ensemble filter from each dataset created by GMM clustering to estimate training dataset quality using the proportion of retained data. The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903. These demonstrate mimic score’s effectiveness for dataset quality estimation.
5.2 Data Selection in Large-scale Web Datasets
Setups.
Next, we test Grad-Mimic in a more challenging setting using million-scale web-crawled datasets. We use the small-scale and medium-scale DataComp datasets [17], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken. This means that our results might not be comparable to previous approaches on the DataComp Leaderboard. See here for details., respectively, to train CLIP model from scratch [2]. In these datasets, noise is naturally presented in the data. We follow the training setup from DataComp [49, 17] and use publicly available pretrained CLIP model’s weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K. This reference model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us. It serves as a proxy for the ideal reference point for scenarios where training such powerful models is infeasible. We use the final MLP layer in the text and image encoders respectively as our target to mimic. We evaluate the performance of resulting model using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks. Other details can be found in Appendix B.2.
In Stage 2, we use top-$k$ percent identification to create the final filter, subsequently training a second CLIP model on the curated dataset to validate resulting filter’s effectiveness.